Please find the value of the expression 4P4
Answer by praseenakos@yahoo.com(417)

Question:

Please find the value of the expression 4P4

Answer:
4P4 = 4* 3 * 2 * 1 = 24

If your question is 4P3.
Then 4P3 = 4 * 3 * 2 = 24


Similarly... 4P2 = 4 * 3 = 12 and

4P1 = 4

Hope you understood..
Regards.

Praseena.